# Isolated zero mode in a quantum computer from a duality twist

### v4

@@ Abstract
Investigating the interplay of dualities, generalized symmetries, and topological defects beyond theoretical models is an important challenge in condensed matter physics and quantum materials. A simple model exhibiting this physics is the transverse-field Ising model, which can host atopological defect that performs the Kramers-Wannier duality transformation. When acting on one point in space, this duality defect imposes the duality twisted boundary condition and binds a single zero mode. This zero mode is unusual as it lacks a localized partner in the same $Z_{2}$ sector and has an infinite lifetime, even in finite systems. Using Floquet driving of a closed Ising chain with a duality defect, we generate this zero mode in a digital quantum computer. We detect the mode by measuring its associated persistent autocorrelation function using an efficient sampling protocol and a compound strategy for error mitigation. We also show that the zero mode resides at the domain wall between two regions related by a Kramers-Wannier duality transformation. Finally, we highlight the robustness of the isolated zero mode to integrability- and symmetry-breaking perturbations. Our findings provide a method for exploring exotic topological defects, associated with noninvertible generalized symmetries, in digitized quantum devices.
I Introduction
@@
The Kramers-Wannier duality transformation of the transverse-field Ising model (TFIM) is one of the most well-known dualities [17]. The TFIM can support different topological defects, including the familiar unitary spin-flip defect, which captures the $Z_{2}$ symmetry of the model, and the duality defect [7, 9, 18] studied here. The duality defect is one of the simplest examples of a noninvertible transformation as it projects out states of a given $Z_{2}$ symmetry [7]. A duality defect acting at a particular point in space imposes a special “twisted” boundary condition [19, 5, 20, 21, 22, 23, 24, 25]. One of the consequences of the duality defect is the emergence of a zero mode, bound to the defect, with distinct characteristics from more well-known edge zero modes occurring in the TFIM with open boundary conditions [7, 18].
It is well-established that the TFIM with open boundary conditions can host Majorana zero modes [26, 27] in the topological phase, both in equilibrium and in its Floquet versions [28, 29]. These modes appear in pairs localized at the two edges of the system, and have a lifetime which grows exponentially with the system size because they hybridize leading to an energy splitting that is exponentially small in the system size. On the other hand, the zero mode appearing due to a duality twist defect is isolated, lacking a localized partner in the same $Z_{2}$ sector. It thus has an infinite lifetime, even in finite systems. In addition, this isolated mode exists for all couplings of the static and Floquet versions of the TFIM, and separates the chain into two parts whose couplings are related by the Kramers-Wannier duality transformation [7, 30].
Digitized quantum devices have provided a new playground for realizing topological physics [31, 32, 33, 34, 35, 36, 37, 38]. In particular, Floquet versions of the TFIM with open boundary conditions have been studied on current noisy quantum devices [39, 37]. In this paper, we probe the Floquet version of the TFIM with a duality twist defect. We implement the duality twist defect in a closed chain of 20 qubits on the ibmq_kolkata device and observe signatures of an isolated zero mode. By obtaining the autocorrelation functions of several operators in the quantum device, we demonstrate that the zero mode, with a theoretically infinite lifetime [18, 30], is persistent. We also confirm that weak perturbations, which break both the integrability and the $Z_{2}$ symmetry of the system, do not destabilize the zero mode in finite-length chains.
@@
The paper is organized as follows. In Section II, we introduce the model and define the autocorrelation function utilized for detecting the zero mode. Sections III, IV, and V are focused on results obtained on the IBM quantum computer. In particular, Sec. III discusses the localized zero mode for the integrable model, Sec. IV demonstrates the topological nature of the duality twist and its relation to the Kramers-Wannier duality transformation, Sec. V explores the effects of integrability- and symmetry breaking perturbations, and Sec. VI summarizes our findings. Additional details are provided in three appendices.
II Unpaired zero mode in the Floquet-TFIM with a duality twist
@@
Duality-twisted boundary conditions, e.g., on the link between sites $r,r-1$, are implemented by changing $\sigma_{r-1}^{z}\sigma_{r}^{z}$ to $\sigma_{r-1}^{z}\sigma_{r}^{x}$, and removing the transverse field $g\sigma_{r}^{x}$ [7], while still preserving periodic boundary conditions with operator $L$ identified with operator $L+j$. The duality twisted Hamiltonian then becomes
@@
$H_{zz}=J\sum_{j\neq r-1}\sigma_{j}^{z}\sigma_{j+1}^{z},\,H_{zx}=J\sigma_{r-1}^{z}\sigma_{r}^{x},\,H_{x}=g\sum_{j\neq r}\sigma_{j}^{x}.$
Note that with the periodic boundary conditions, in the absence of a twist, the $Z_{2}$ symmetry is ${\mathcal{D}}_{\psi}=\prod_{j=0}^{L-1}\sigma^{x}_{j}$. With the twist, the $Z_{2}$ symmetry changes to $\Omega_{r}=i\sigma^{z}_{r}\mathcal{D}_{\psi}$. A duality twist can also be imposed by considering an infinite chain, and gauging the $Z_{2}$ symmetry for one half of the chain [40, 41, 42, 43, 44, 45] or by performing a Kramers-Wannier transformation on one half of the chain [46]. One can also derive the above duality twisted Hamiltonian by using techniques from the Ising fusion category [46].
@@
$\gamma_{2(r+j)}=\sigma^{z}_{j+r+1}\prod_{k=r}^{j+r}\sigma^{x}_{k},\,\,j=0,1,\ldots L-2,$ (2a)
$\gamma_{2(r+j)+1}=-\sigma^{y}_{j+r+1}\prod_{k=r}^{j+r}\sigma^{x}_{k},\,\,j=0,1,\ldots L-2,$ (2b)
$\gamma_{2r+2L-2}=\sigma^{z}_{r},\,\,\gamma_{2r+2L-1}=-\sigma^{y}_{r},$ (2c)
$\sigma_{L+j}^{\alpha}=\sigma_{j}^{\alpha}.$ (2d)
Above, the last equation imposes periodic boundary conditions. With the twist, the $Z_{2}$ symmetry in terms of the Majorana fermions is
$\Omega_{r}=i\sigma^{z}_{r}\mathcal{D}_{\psi}=\gamma_{2r+2L-1}\biggl{[}\left(-i\gamma_{2r}\gamma_{2r+1}\right)\left(-i\gamma_{2r+2}\gamma_{2r+3}\right)\ldots\left(-i\gamma_{2r+2L-4}\gamma_{2r+2L-3}\right)\biggr{]}.$ (3)
Note that $\Omega_{r}$ is a product of $12L-$ 1 Majorana fermions as it does not contain the $\gamma_{2r+2L-2}$ Majorana fermion. Since it is a string containing odd number of Majorana fermions, we have the following relations
$\Omega_{r}\gamma_{2r+2L-2}=-\gamma_{2r+2L-2}\Omega_{r},$ (4)
[…]
@@
$H_{zz}=-iJ\sum_{j=0\ldots L-2,j\neq L-1}\gamma_{2(j+r)-1}\gamma_{2(j+r)},$ (7)
$H_{x}=-ig\sum_{j=0\ldots L-2,j\neq L-1}\gamma_{2(r+j)}\gamma_{2(r+j)+1},$ (8)
$H_{zx}=-iJ\Omega_{r}\gamma_{2(r+L)-3}\gamma_{2(r+L)-1}.$ (9)
@@
Even though there are $2L$ Majorana fermions in the Hilbert space, $H_{t}$ explicitly contains $12L-$ 1 Majorana fermions as $\gamma_{2(r+L)-2}$ does not enter in $H_{t}$. Importantly, although $\gamma_{2r+2L-2}$ does not appear in the Hamiltonian, it is not a zero mode as it does not commute with $\Omega_{r}$ and consequently the Hamiltonian. The string operator appears to make the system interacting, but the problem is still solvable because the $12L-$ 1 Majorana fermions entering in $H_{t}$ commute with $\Omega_{r}$. Thus one may set $\Omega_{r}=\pm 1$, after which the Hamiltonian is a bilinear in the Majorana fermions. Thus $H_{t}$, has eigenmodes obtained by diagonalizing a $(2L-1)\times(2L-1)$ matrix for each $Z_{2}$ sector. Since the matrix dimension is odd in each $Z_{2}$ sector, this has to yield an odd number of zero modes in each $Z_{2}$ sector.
@@
For the parameters above, a given $Z_{2}$ sector hosts a zero mode that is localized at $r$, and has a localization length $\xi$, where $\exp[1/\xi]\approx\text{max}(g,J)/\text{min}(g,J)$ [18]. There is nothing special about the sites $r,r-1$, and a local unitary transformation can move the twist defect from one pair of sites to the neighboring pair of sites [47, 7, 30] without moving the zero mode [30]. In particular, after a JW transformation this zero mode is a Majorana mode, with the branch-cut or string associated with the zero mode terminating at the duality defect [7]. The local unitary transformation just modifies the length of this branch cut as it does not move the zero mode while moving the site of the duality defect.
@@
Some approaches in the literature (see [48, 49] and appendix of [7]) have also considered a related model that is more physical for fermions, where the string operator is simply dropped to obtain the new model. In this case, there are indeed two zero modes with the second zero mode being $\gamma_{2(r+L)-2}$ which now commutes with the Hamiltonian (as the string operator does not explicitly enter in the Hamiltonian). However, this is a sleight of hand, which is strictly not valid if one is discussing the physics of the original spin chain. In other words, the model with two zero modes where one of them is $\gamma_{2(r+L)-2}$, is a different model. For the spin chain, a proper treatment always leads to the string operator $\Omega_{r}$, and there is no good reason to drop it. In this case $\gamma_{2(r+L)-2}$ is not a zero mode as it does not commute with $H_{t}$ due to the string $\Omega_{r}$.
@@
In summary, the spin chain, has one zero mode in each $Z_{2}$ sector, thus when looking at both even and odd $Z_{2}$ sectors, $\Omega_{r}=\pm 1$, there are two zero modes. In what follows, we will perform simulations that do not mix $Z_{2}$ sectors. In this case, only a single zero mode will manifest. The above method of solving the problem in each $Z_{2}$ sector separately appears in other contexts such as the periodic chain. For this case, the JW transformation to fermions gives a nonlocal term on one bond where the nonlocal term corresponds to fermion parity that serves as the $Z_{2}$ symmetry of the model. Thus, in each symmetry sector, the nonlocal term is replaced by its ±$\pm 1$ eigenvalue, making the model local.
@@
As far as symmetries go, much of what was discussed regarding $H_{t}$ is also valid for $U$. In particular, in a given $Z_{2}$ sector $\Omega_{r}=\pm 1$, $U$ is exactly solvable as it involves exponentials of fermion bilinears after a JW transformation. In addition, only $12L-$ 1 Majorana fermions appear in $U$, so that we need to diagonalize a $(2L-1)\times(2L-1)$ dimensional matrix. Since the orthogonal matrix that evolves the $12L-$ 1 Majoranas only performs proper rotations, its determinant has to be $1$. This forces an odd number of eigenvalues to be equal to $1$, where the corresponding eigenvectors are the zero modes.
The analytic expression for the zero mode for the above Floquet set-up, i.e., with a duality-twist, was derived in [18, 30]. Similar to edge modes with open boundary conditions, the zero mode with a duality twist is a linear superposition of Majoranas all along the chain. For the zero mode with the duality twist, the largest weight in the superposition is with the Majorana in the vicinity of site $r$. In contrast, the zero mode pairs with open boundary conditions have their largest weight at the boundaries. In addition, one may introduce a single domain wall in the duality-twisted Floquet set-up by making the couplings $g,J$ inhomogeneous [18]. In this case, the zero mode resides at the domain wall. As we shall discuss further, a series of local unitary transformations can relate the homogeneous Floquet unitary of (10) to one where a single domain wall is present [30].
The Majorana zero mode with a duality twist has important differences with the Majorana pairs that appear for open boundary conditions. While the latter occurs only in the topological phase $J>g$, the Majorana zero mode with the duality twist occurs for all values of $g,J$, including $g>J,g<J$ and even disordered ones. Only its localization length is determined by the strength of the disorder and average values of $g$ and $J$. Since the zero mode with the duality twist lacks a partner, it never decays in time, i.e., it has an infinite lifetime rather than a lifetime that is exponentially large in the system size. The mathematical reason for this is that the zero mode with the duality twist is a symmetry of the system as it commutes both with the generator of time evolution $U$, as well as the $Z_{2}$ symmetry $\Omega_{r}$. In contrast, the Majorana pairs with open boundary conditions anticommute with the associated $Z_{2}$ symmetry. We emphasize that a single zero mode in a given $Z_{2}$ sector for the duality twisted model follows from a simple counting argument as summarized above.
@@
Figure 1: Implementation of the unitary (10) for a 4-qubit system in a quantum computer. The duality twist corresponds to modifying the Floquet TFIM by removing the transverse field on qubit $3$ and applying a twist exchange interaction $\sigma^{z}_{2}\sigma^{x}_{3}$ between qubits $2$ and $3$. Periodic boundary conditions have been imposed with a $\sigma^{z}_{3}\sigma^{z}_{0}$ exchange interaction.
Figure 2: Signature of the zero mode at the defect site for the duality-twisted Floquet-TFIM. The label $q_{i}$ denotes the measurement at the $i$th qubit. Panel (a) shows the nondecaying autocorrelation for $\sigma_{y}$ at the defect site $r=L-1$, which overlaps with the zero mode. All other autocorrelations both at the twist defect $r=L-1$ [panel (b), (c)] and at a bulk site $j=-1+L/2$ [panel (d), (e), (f)], far away from the twist defect, decay with stroboscopic time $n$. The chosen bulk qubit behaves qualitatively similarly to other bulk qubits. We present data from ibmq_kolkata device for $L=20$, the longest physical loop of qubits available in the device. There is good qualitative agreement between noiseless simulation and measurements on the noisy quantum device. The noiseless simulations were also performed for a smaller system of 12 qubits. The agreement between the two system sizes $L=12,20$ indicates that we are capturing the behavior in the thermodynamic limit. Date of demonstration at ibmq_kolkata device is 18th April, 2023.
III Detection of the unpaired zero mode in the IBM quantum device
We have chosen the values of the coupling constants to be $J=0.75\pi$ and $g=0.2\pi$ for all the computations in this paper, although similar results can be obtained for other coupling constants. To create and detect the zero mode in the quantum computer, we apply a quantum circuit that generates the Floquet dynamics by the repeated applications of a circuit block corresponding to the unitary operator in (10). A simple 4-qubit version of this block, corresponding to one Trotter step, is shown in Fig. 1. We compute the autocorrelation (11) for $\mathcal{O}=\sigma^{x},\sigma^{y}$, and $\sigma^{z}$ of a 20-qubit system. We perform this calculation on the 27-qubit ibmq_kolkata device with a 20-qubit closed loop. Thus, the device geometry eliminates the need for using swap gates to implement periodic boundary conditions. We used a compound error mitigation method as discussed in Appendix B and described in Refs. [50, 38].

### v6

@@ Abstract
Investigating the interplay of dualities, generalized symmetries, and topological defects beyond theoretical models is an important challenge in condensed matter physics and quantum materials. A simple model exhibiting this physics is the transverse-field Ising model, which can host a topological defect that performs the Kramers-Wannier duality transformation. When acting on one point in space, this duality defect imposes the duality twisted boundary condition and binds a single zero mode. This zero mode is unusual as it lacks a localized partner in the same $\mathbb{Z}_{2}$ sector and has an infinite lifetime, even in finite systems. Using Floquet driving of a closed Ising chain with a duality defect, we generate this zero mode in a digital quantum computer. We detect the mode by measuring its associated persistent autocorrelation function using an efficient sampling protocol and a compound strategy for error mitigation. We also show that the zero mode resides at the domain wall between two regions related by a Kramers-Wannier duality transformation. Finally, we highlight the robustness of the isolated zero mode to integrability- and symmetry-breaking perturbations. Our findings provide a method for exploring exotic topological defects, associated with noninvertible generalized symmetries, in digitized quantum devices.
1 Introduction
@@
The Kramers-Wannier (KW) duality transformation of the transverse-field Ising model (TFIM) is one of the most well-known dualities [17]. The TFIM can support different topological defects, including the familiar unitary spin-flip defect, which captures the $\mathbb{Z}_{2}$ symmetry of the model, and the duality defect [7, 9, 18] studied here.The duality defect is one of the simplest examples of a noninvertible transformation as it projects out states of a given $\mathbb{Z}_{2}$ symmetry [7]. Recently, the integrability of the critical TFIM was studied in Ref. [19] (see also Ref. [20] for a wave-function perspective) and many conserved charges were found, including charges associated with the KW duality and other noninvertible symmetries. A duality defect acting at a particular point in space imposes a special “twisted” boundary condition [21, 5, 22, 23, 24, 25, 26, 27]. One of the consequences of the duality defect is the emergence of a zero mode, bound to the defect, with distinct characteristics from more well-known edge zero modes occurring in the TFIM with open boundary conditions [7, 18].
It is well-established that the TFIM with open boundary conditions can host Majorana zero modes [28, 29] in the topological phase, both in equilibrium and in its Floquet versions [30, 31]. These modes appear in pairs localized at the two edges of the system, and have a lifetime which grows exponentially with the system size because they hybridize leading to an energy splitting that is exponentially small in the system size. On the other hand, the zero mode appearing due to a duality twist defect is isolated, lacking a localized partner in the same $\mathbb{Z}_{2}$ sector. It thus has an infinite lifetime, even in finite systems. In addition, this isolated mode exists for all couplings of the static and Floquet versions of the TFIM, and separates the chain into two parts whose couplings are related by the Kramers-Wannier duality transformation [7, 32].
Digitized quantum devices have provided a new playground for realizing topological physics [33, 34, 35, 36, 37, 38, 39, 40]. In particular, Floquet versions of the TFIM with open boundary conditions have been studied on current noisy quantum devices [41, 39]. In this paper, we probe the Floquet version of the TFIM with a duality twist defect. We implement the duality twist defect in a closed chain of 20 qubits on the ibmq_kolkata device and observe signatures of an isolated zero mode. By obtaining the autocorrelation functions of several operators in the quantum device, we demonstrate that the zero mode, with a theoretically infinite lifetime [18, 32], is persistent. We also confirm that weak perturbations, which break both the integrability and the $\mathbb{Z}_{2}$ symmetry of the system, do not destabilize the zero mode in finite-length chains.
@@
The paper is organized as follows. In Section 2, we introduce the model and define the autocorrelation function utilized for detecting the zero mode. Sections 3, 4, and 5 are focused on results obtained on the IBM quantum computer. In particular, Sec. 3 discusses the localized zero mode for the integrable model, Sec. 4 demonstrates the topological nature of the duality twist and its relation to the Kramers-Wannier duality transformation, Sec. 5 explores the effects of integrability- and symmetry-breaking perturbations, and Sec. 6 summarizes our findings. Additional details are provided in three appendices.
2 Unpaired zero mode in the Floquet-TFIM with a duality twist
@@
Duality-twisted boundary conditions, e.g., on the link between sites r,r1r,r-1, are implemented by changing $\sigma_{r-1}^{z}\sigma_{r}^{z}$ to $\sigma_{r-1}^{z}\sigma_{r}^{x}$, and removing the transverse field $g\sigma_{r}^{x}$ [7], while still preserving periodic boundary conditions with operator $L$ identified with operator $L+j$. The duality twisted Hamiltonian then becomes
@@
$H_{zz}=J\sum_{j\neq r-1}\sigma_{j}^{z}\sigma_{j+1}^{z},\,H_{zx}=J\sigma_{r-1}^{z}\sigma_{r}^{x},$
$H_{x}=g\sum_{j\neq r}\sigma_{j}^{x}.$
Note that with the periodic boundary conditions, in the absence of a twist, the $\mathbb{Z}_{2}$ symmetry is ${\mathcal{D}}_{\psi}=\prod_{j=0}^{L-1}\sigma^{x}_{j}$. With the twist, the $\mathbb{Z}_{2}$ symmetry changes to $\Omega_{r}=i\sigma^{z}_{r}\mathcal{D}_{\psi}$. A duality twist can also be imposed by considering an infinite chain, and gauging the $\mathbb{Z}_{2}$ symmetry for one half of the chain [42, 43, 44, 45, 46, 47] or by performing a Kramers-Wannier transformation on one half of the chain [48]. One can also derive the above duality twisted Hamiltonian by using techniques from the Ising fusion category [48].
@@
$\gamma_{2j}=\sigma^{z}_{j}\prod_{k=r}^{j-1}\sigma^{x}_{k},\,\,$ (2a)
$\gamma_{2j+1}=-\sigma^{y}_{j}\prod_{k=r}^{j-1}\sigma^{x}_{k},\,\,$ (2b)
$\gamma_{2r}=\sigma^{z}_{r},\,\,\gamma_{2r+1}=-\sigma^{y}_{r}.$ (2c)
In the above equations $j\neq r,j=0\ldots r-1,r+1,\ldots L-1$. It is understood that as the Majorana index $j$ increases, the JW string starts from site $r$, reaches L1L-1 and then continues to site $0,1\ldots$. In particular $\gamma_{2r-1}=-\sigma^{y}_{r-1}{\mathcal{D}}_{\psi}\sigma^{x}_{r-1}$. We will use this below when writing the twist term $H_{zx}$ in terms of Majorana fermions.
With the twist, the $\mathbb{Z}_{2}$ symmetry in terms of the Majorana fermions is
$\Omega_{r}=i\sigma^{z}_{r}\mathcal{D}_{\psi}=(-i)^{L-1}\biggl[\gamma_{0}\gamma_{1}\gamma_{2}\gamma_{3}\ldots$
$\gamma_{2r-2}\gamma_{2r-1}\gamma_{2r+1}\gamma_{2r+2}\gamma_{2r+3}\ldots\gamma_{2L-2}\gamma_{2L-1}\biggr].$ (3)
Note that $\Omega_{r}$ is a product of 2L12L-1 Majorana fermions as it does not contain the $\gamma_{2r}$ Majorana fermion. Since it is a string containing odd number of Majorana fermions, we have the following relations
$\Omega_{r}\gamma_{2r}=-\gamma_{2r}\Omega_{r},$ (4)
$\Omega_{r}\gamma_{j}=\gamma_{j}\Omega_{r},\hskip 46.94687ptj\neq 2r$ (5)
@@
$H_{zz}=-iJ\sum_{j=0\ldots L-1,j\neq r}\gamma_{2j-1}\gamma_{2j},$ (6)
$H_{x}=-ig\sum_{j=0\ldots L-1,j\neq r}\gamma_{2j}\gamma_{2j+1},$ (7)
$H_{zx}=-iJ\Omega_{r}\gamma_{2r-1}\gamma_{2r+1}.$ (8)
@@
Even though there are $2L$ Majorana fermions in the Hilbert space, $H_{t}$ explicitly contains 2L12L-1 Majorana fermions as $\gamma_{2r}$ does not enter in $H_{t}$. Importantly, although $\gamma_{2r}$ does not appear in the Hamiltonian, it is not a zero mode as it does not commute with $\Omega_{r}$ and consequently the Hamiltonian. The string operator appears to make the system interacting, but the problem is still solvable because the 2L12L-1 Majorana fermions entering in $H_{t}$ commute with $\Omega_{r}$. Thus one may set $\Omega_{r}=\pm 1$, after which the Hamiltonian is a bilinear in the Majorana fermions. Thus $H_{t}$, has eigenmodes obtained by diagonalizing a $(2L-1)\times(2L-1)$ matrix for each $\mathbb{Z}_{2}$ sector. Since the matrix dimension is odd in each $\mathbb{Z}_{2}$ sector, this has to yield an odd number of zero modes in each $\mathbb{Z}_{2}$ sector.
@@
For the parameters above, a given $\mathbb{Z}_{2}$ sector hosts a zero mode that is localized at $r$, and has a localization length $\xi$, where $\exp[1/\xi]\approx\text{max}(g,J)/\text{min}(g,J)$ [18]. There is nothing special about the sites r,r1r,r-1, and a local unitary transformation can move the twist defect from one pair of sites to the neighboring pair of sites [49, 7, 32] without moving the zero mode [32]. In particular, after a JW transformation this zero mode is a Majorana mode, with the branch-cut or string associated with the zero mode terminating at the duality defect [7]. The local unitary transformation just modifies the length of this branch cut as it does not move the zero mode while moving the site of the duality defect.
@@
Some approaches in the literature (see [50, 51] and appendix of [7]) have also considered a related model that is more physical for fermions, where the string operator is simply dropped to obtain the new model. In this case, there are indeed two zero modes with the second zero mode being $\gamma_{2r}$ which now commutes with the Hamiltonian (as the string operator does not explicitly enter in the Hamiltonian). However, this is a sleight of hand, which is strictly not valid if one is discussing the physics of the original spin chain. In other words, the model with two zero modes where one of them is $\gamma_{2r}$, is a different model. For the spin chain, a proper treatment always leads to the string operator $\Omega_{r}$, and there is no good reason to drop it. In this case $\gamma_{2r}$ is not a zero mode as it does not commute with $H_{t}$ due to the string $\Omega_{r}$.
@@
In summary, the spin chain, has one zero mode in each $\mathbb{Z}_{2}$ sector, thus when looking at both even and odd $\mathbb{Z}_{2}$ sectors, $\Omega_{r}=\pm 1$, there are two zero modes. In what follows, we will perform simulations that do not mix $\mathbb{Z}_{2}$ sectors. In this case, only a single zero mode will manifest. The above method of solving the problem in each $\mathbb{Z}_{2}$ sector separately appears in other contexts such as the periodic chain. For this case, the JW transformation to fermions gives a nonlocal term on one bond where the nonlocal term corresponds to fermion parity that serves as the $\mathbb{Z}_{2}$ symmetry of the model. Thus, in each symmetry sector, the nonlocal term is replaced by its $\pm 1$ eigenvalue, making the model local.
@@
As far as symmetries go, much of what was discussed regarding $H_{t}$ is also valid for $U$. In particular, in a given $\mathbb{Z}_{2}$ sector $\Omega_{r}=\pm 1$, $U$ is exactly solvable as it involves exponentials of fermion bilinears after a JW transformation. In addition, only 2L12L-1 Majorana fermions appear in $U$, so that we need to diagonalize a $(2L-1)\times(2L-1)$ dimensional matrix. Since the orthogonal matrix that evolves the 2L12L-1 Majoranas only performs proper rotations, its determinant has to be $1$. This forces an odd number of eigenvalues to be equal to $1$, where the corresponding eigenvectors are the zero modes.
The analytic expression for the zero mode for the above Floquet set-up, i.e., with a duality-twist, was derived in [18, 32]. Similar to edge modes with open boundary conditions, the zero mode with a duality twist is a linear superposition of Majoranas all along the chain. For the zero mode with the duality twist, the largest weight in the superposition is with the Majorana in the vicinity of twist. In contrast, the zero mode pairs with open boundary conditions have their largest weight at the boundaries. In addition, one may introduce a single domain wall in the duality-twisted Floquet set-up by making the couplings $g,J$ inhomogeneous [18]. In this case, the zero mode resides at the domain wall. As we shall discuss further, a series of local unitary transformations can relate the homogeneous Floquet unitary of (9) to one where a single domain wall is present [32].
The Majorana zero mode with a duality twist has important differences with the Majorana pairs that appear for open boundary conditions. While the latter occurs only in the topological phase $J>g$, the Majorana zero mode with the duality twist occurs for all values of $g,J$, including $g>J,g<J$ and even disordered ones. Only its localization length is determined by the strength of the disorder and average values of $g$ and $J$. Since the zero mode with the duality twist lacks a partner, it never decays in time, i.e., it has an infinite lifetime rather than a lifetime that is exponentially large in the system size. The mathematical reason for this is that the zero mode with the duality twist is a symmetry of the system as it commutes both with the generator of time evolution $U$, as well as the $\mathbb{Z}_{2}$ symmetry $\Omega_{r}$. In contrast, the Majorana pairs with open boundary conditions anticommute with the associated $\mathbb{Z}_{2}$ symmetry. We emphasize that a single zero mode in a given $\mathbb{Z}_{2}$ sector for the duality twisted model follows from a simple counting argument as summarized above.
@@
Figure 1: Implementation of the unitary (9) for a 4-qubit system in a quantum computer. The duality twist corresponds to modifying the Floquet TFIM by removing the transverse field on qubit 33 and applying a twist exchange interaction $\sigma^{z}_{2}\sigma^{x}_{3}$ between qubits 22 and $3$. Periodic boundary conditions have been imposed with a $\sigma^{z}_{3}\sigma^{z}_{0}$ exchange interaction.
Figure 2: Signature of the zero mode at the defect site for the duality-twisted Floquet-TFIM. The label $q_{i}$ denotes the measurement at the $i$th qubit. Panel (a) shows the nondecaying autocorrelation for $\sigma^{y}_{r}$ at the defect site r=L1r=L-1, which overlaps with the zero mode. All other autocorrelations both at the twist defect r=L1r=L-1 [panel (b), (c)] and at a bulk site j=1+L/2j=-1+L/2 [panel (d), (e), (f)], far away from the twist defect, decay with stroboscopic time $n$. The chosen bulk qubit behaves qualitatively similarly to other bulk qubits. We present data from ibmq_kolkata device for $L=20$, the longest physical loop of qubits available in the device. There is good qualitative agreement between noiseless simulation and measurements on the noisy quantum device. The noiseless simulations were also performed for a smaller system of 12 qubits. The agreement between the two system sizes $L=12,20$ indicates that we are capturing the behavior in the thermodynamic limit. Date of demonstration at ibmq_kolkata device is 18th April, 2023.
3 Detection of the unpaired zero mode in the IBM quantum device
We have chosen the values of the coupling constants to be $J=0.75\pi$ and $g=0.2\pi$ for all the computations in this paper, although similar results can be obtained for other coupling constants. To create and detect the zero mode in the quantum computer, we apply a quantum circuit that generates the Floquet dynamics by the repeated applications of a circuit block corresponding to the unitary operator in (9). A simple 4-qubit version of this block, corresponding to one Trotter step, is shown in Fig. 1. We compute the autocorrelation (10) for $\mathcal{O}=\sigma^{x},\sigma^{y}$, and $\sigma^{z}$ of a 20-qubit system. We perform this calculation on the 27-qubit ibmq_kolkata device with a 20-qubit closed loop. Thus, the device geometry eliminates the need for using swap gates to implement periodic boundary conditions. We used a compound error mitigation method as discussed in Appendix C and described in Refs. [39, 40].